Circular No. 8628

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2005io, 2005ip, 2005iq Three apparent supernovae have been reported from unfiltered CCD

images: 2005io independently by E. Lee and W. Li (via LOSS/KAIT; cf. IAUC 8625; the tabulated data below for 2005io are from KAIT) and by T. Boles (Coddenham, England, 0.35-m reflector); 2005ip by Boles; and 2005iq by H. Khandrika and W. Li (LOSS/KAIT). Additional magnitudes for 2005io for SN 2005io in UGC 3361: 1989 Dec. 3, [21.0 (Digitized Sky Survey red plate; via Boles); 1992 Nov. 24, [21.5 (DSS, blue); 2005 Mar. 14 UT, [19.5 (Boles); Oct. 10, [19.5 (Boles); 28.46, [19.5 (KAIT); Nov. 4.007, 18.6 (Boles; position end figures 24^{s}.62, 06′′.4). Additional magnitudes for 2005ip in NGC 2906: 1994 Feb. 9, [20.5 (DSS, blue, via Boles); 1999 Feb. 20, [21.0 (DSS, red); 2004 Dec. 6 and 2005 Jan. 21, [19.5 (Boles); Nov. 5.261, 14.6 (C. Colesanti, C. Jacques, E. Pimentel, and T. Napoleao, Belo Horizonte, Brazil); 5.682, 14.8 (K. Itagaki, Yamagata, Japan; via S. Nakano, Sumoto; position end figures 06^{s}.40, 44′′.2). Additional KAIT magnitudes for 2005iq in MCG –03-1-8: Oct. 9.22, [19.0; Nov. 6.21, 17.0.

SN  2005 UT  α_{2000}  δ_{2000}  Mag.  Offset
2005io Nov. 3.47 5^{h}50^{m}24.^{s}64 +49^{o}43′06′′.5 18.5 6′′.1 W, 17′′.6 N
2005ip Nov. 5.163 9 32 06.42 + 8 26 44.4 15.5 2′′.8 E, 14′′.2 N
2005iq Nov. 5.20 23 58 32.50 −18 42 33.0 17.2 8′′.6 E, 17′′.7 N

A. V. Filippenko and R. J. Foley, University of California, Berkeley, report that inspection of CCD spectra (range 330–920 nm), obtained on Nov. 5 UT with the Keck I 10-m telescope (+ LRIS), shows that SN 2005io is of type II, shortly after the explosion. The broad hydrogen Balmer lines exhibit P-Cyg profiles, but in the case of Hα the emission component greatly dominates over the absorption component. The continuum is very blue, and He I 587-nm absorption is present.

M. Modjaz, R. Kirshner, and P. Challis, Harvard-Smithsonian Center for Astrophysics, report that a spectrogram (range 350–740 nm) of 2005ip, obtained by M. Calkins on Nov. 6.50 UT with the Mt. Hopkins 1.5-m telescope (+ FAST), shows it to be a type-II supernova, probably within a few weeks past explosion. The spectrum consists of a blue continuum (indicating a young age), the P-Cyg line of Hα, and other absorption troughs. Adopting a recession velocity of 2440 km/s for the host galaxy (from Falco et al. 1999, PASP 111, 438), the expansion velocity derived from the minimum of the Hα line is 15400 km/s. An absorption trough observed at ∼ 425.0 nm seems to be too blue to be identified with Hβ.

2005 November 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green